# Who’s the GOAT? Sports Rankings and Data-Driven Random Walks on the Symmetric Group

### v1

@@ 4.3. Rankings from the WTA
In Figure 7 we plot the top portion of the Hasse diagram of the poset over “top 5 players,” where Graf and S. Williams are again incomparable maximal elements. Note that while Graf is better than Navratilova, S. Williams and Navratilova are incomparable. Moreover, note that unlike the case in Figure 6, as top 5 players, both Clijsters and Sharapova are better than Sabalenka. This is a similar effect to that seen with Alcaraz in Section 4.2: Sabalenka turned professional in 2015 and had little overlap in competition with Sharapova, who retired in 2020, and Clijsters, who retired in 2012 except for a brief comeback between 2020 and 2022. In particular, her prevalence within the top 3 is significant independently of her short career at the time of writing. However, when the cutoff is relaxed to top 5, her short career-to-date puts her at a disadvantage compared to long-established players with significant career prevalence in the top 5, such as Clijsters and Sharapova. Qualitatively, we observe the inclusion of more players near the top of this Hasse diagram compared to the previous cases, and that chains of players that were comparable as top 3 players become anti-chains and close to maximal as top 5 players (e.g., Lindsay Davenport and Seles).
In Figure 8 we plot the top portion of the Hasse diagram of the poset over “top 10 players,” where now the three of Graf, Navratilova, and S. Williams are incomparable maximal elements. In other words, while Navratilova is close to but not quite the GOAT as a top 3 or top 5 player, she is conclusively one of the GOATs as a top 10 player. As a top 10 player, Navratilova is better than Seles. However, Graf and S. Williams cannot be conclusively compared to Seles (even though they fare slightly better than her in terms of average rank).
@@
Lastly, in Figure 9 we plot the top portion of the Hasse diagram of the poset over “top 20 players,” where now all six of Graf, Navratilova, Seles, S. Williams, V. Williams, and Wozniacki are incomparable maximal elements. Qualitatively, we observe a significant and across the board increase in the number of players near the top of the Hasse diagram. This indicates that, in the WTA, there are many more good top 20 players than there are good top 3, 5, or 10 players.
@@ 4.4. Comparing the Associations
Figures 5 and 9 are qualitatively different in that, as top 20 players, the number of players close to maximal is much larger for the WTA than it is for the ATP. Comparing the results from Sections 4.2 and 4.3, this difference is not nearly as apparent for the posets over top 3, 5, or 10 players. Here, we further investigate this difference from a more quantitative point of view. In particular, we ask questions of the form: to what extent is being “a good top $\kappa$ player” a predictor of being “a good top $\kappa^{\prime}$ player,” where $\kappa^{\prime}\leq\kappa$? We formalize the notion of “a good top $\kappa$ player” by average rank (from the linear extensions) for the poset cutoff $\kappa$. Intuitively, the smaller the difference between $\kappa$ and $\kappa^{\prime}$, the stronger we expect the correlation to be.
In Figure 10, we consider the set of players that appear among the first $25$ by average rank for all cutoffs $\kappa=3,5,10,20$. This amounts to $15$ players in both the ATP and in the WTA. Then, for each of these sets of $15$ players, we plot a matrix of scatter plots of average rank and varying cutoffs. In this way, each point represents a player, and its coordinates correspond to the player’s average rank for a combination of cutoffs. For each scatter plot, we also fit a linear model and report its coefficient of determination $r^{2}$.
@@
From Figure 10 we note that, in both associations, being a good top 5 player is a reasonable predictor of being a good top 3 player, as is being a good top 10 player a reasonable predictor for being a good top 5 player. In both cases, being a good top 10 player is still correlated with being a good top 3 player, although much less so than with being a good top 5 player. However, when we consider being a good top 20 player as a predictor, we observe a major difference between the two associations. In the ATP, being a good top 20 player as a predictor is correlated with being a good top 10, 5, and 3 player (with decreasing correlation in that order). Conversely, in the WTA, being a good top 20 player is only mildly correlated with being a good top 10 player, and it is nearly uncorrelated with being a good top 5 or top 3 player. This quantitative finding highlights a major difference in terms of career and dominance longevity for top players across the two associations. While we do not delve into causality in this work, we believe it would be practically important to further study the reasons and broader implications of this phenomenon.
@@ 5. Conclusion
In this work, we proposed a mathematical framework to compare players, and in particular players spanning across different time periods in sports history. We aggregated cutoffs of historical ranking data into a single poset over top players by i) sampling from a data-driven random walk over the symmetric group and ii) adopting a notion of stochastic dominance. Ultimately, some pairs of players can be conclusively compared while others cannot. This approach departs from existing methods in that it relies on ranking data (rather than match-level data, which may or may not exist/be available) and formally comes to terms with the possibility that some comparisons are “too close to call.” The two steps in our framework leave ample room for further modeling choices, namely by adapting the construction of (1) and (4) as needed for the particular application in mind. These applications also motivate the analytical study of data-driven sampling of permutations.
We implemented our methods using data from both the ATP and the WTA to find the GOATs in the respective categories. Our experimental results suggest that Nadal, Federer, and Djokovic are the ones that tend to come ahead as GOATs of the ATP under most circumstances, whereas Graf and S. Williams are the ones that do so in the case of the WTA. Through our experiments, we also noted major differences in terms of career and dominance longevity for top players across the two associations.
@@ Acknowledgements
Appendix
The full Hasse diagrams from the experiments in Section 4 are shown in Figure 11.
(a) Rankings from the ATP.
(b) Rankings from the WTA.

### v2

@@ 4.3. Rankings from the WTA
In Figure 9 we plot the top portion of the Hasse diagram of the poset over “top 5 players,” where Graf and S. Williams are again incomparable maximal elements. Note that while Graf is better than Navratilova, S. Williams and Navratilova are incomparable. Moreover, note that unlike the case in Figure 8, as top 5 players, both Clijsters and Sharapova are better than Sabalenka. This is a similar effect to that seen with Alcaraz in Section 4.2: Sabalenka turned professional in 2015 and had little overlap in competition with Sharapova, who retired in 2020, or with Clijsters, who retired in 2012 except for a brief comeback between 2020 and 2022. In particular, her prevalence within the top 3 is significant independently of her short career at the time of writing. However, when the cutoff is relaxed to top 5, her short career-to-date puts her at a disadvantage compared to long-established players with significant career prevalence in the top 5, such as Clijsters and Sharapova. Qualitatively, we observe the inclusion of more players near the top of this Hasse diagram compared to the previous cases, and that chains of players that were comparable as top 3 players become anti-chains and close to maximal as top 5 players (e.g., Lindsay Davenport and Seles).
In Figure 10 we plot the top portion of the Hasse diagram of the poset over “top 10 players,” where now the three of Graf, Navratilova, and S. Williams are incomparable maximal elements. In other words, while Navratilova is close to but not quite the GOAT as a top 3 or top 5 player, she is conclusively one of the GOATs as a top 10 player. As a top 10 player, both Graf and Navratilova are better than Seles. However, S. Williams cannot be conclusively compared to Seles even though she fares better than her in terms of average rank.
@@
Lastly, in Figure 11 we plot the top portion of the Hasse diagram of the poset over “top 20 players,” where now all five of Graf, Navratilova, S. Williams, V. Williams, and Wozniacki are incomparable maximal elements. Qualitatively, we observe a significant and across the board increment in the number of players near the top of the Hasse diagram. This indicates that, in the WTA, there are many more good top 20 players than there are good top 3, 5, or 10 players.
@@ 4.4. Comparing the Associations
Figures 7 and 11 are qualitatively different in that, as top 20 players, the number of players close to maximal is much larger for the WTA than it is for the ATP. Comparing the results from Sections 4.2 and 4.3, this difference is not nearly as apparent for the posets over top 3, 5, or 10 players. Here, we further investigate this difference from a more quantitative point of view. In particular, we ask questions of the form: to what extent is being “a good top $\kappa$ player” a predictor of being “a good top $\kappa^{\prime}$ player,” where $\kappa^{\prime}\leq\kappa$? We formalize the notion of “a good top $\kappa$ player” by average rank (from the linear extensions) for the poset cutoff $\kappa$. Intuitively, the smaller the difference between $\kappa$ and $\kappa^{\prime}$, the stronger we expect the correlation to be.
In Figure 12, we consider the set of players that appear among the first $50$ by average rank for all cutoffs $\kappa=3,5,10,20$. This amounts to $35$ players in the WTA and $32$ the ATP. Then, for each of these sets of players, we plot a matrix of scatter plots of average rank and varying cutoffs. In this way, each point represents a player, and its coordinates correspond to the player’s average rank for a combination of cutoffs. For each scatter plot, we also fit a linear model and report its coefficient of determination $r^{2}$.
@@
From Figure 12 we note that, in both associations, being a good top 5 player is a reasonable predictor of being a good top 3 player, as is being a good top 10 player a reasonable predictor for being a good top 5 player. In both cases, being a good top 10 player is still correlated with being a good top 3 player, although much less so than with being a good top 5 player. However, when we consider being a good top 20 player as a predictor, we observe a major difference between the two associations. In the ATP, being a good top 20 player is correlated with being a good top 10, 5, and 3 player (with decreasing correlation in that order). Conversely, in the WTA, being a good top 20 player is only mildly correlated with being a good top 10 player, and it is nearly uncorrelated with being a good top 5 or top 3 player. In general, all correlations are weaker in the WTA than they are in the ATP. This quantitative finding highlights a major difference in terms of career and dominance longevity for top players across the two associations. While we do not delve into causality in this work, we believe it would be practically important to further study the reasons and broader implications of this phenomenon.
@@ 5. Conclusion
In this work, we proposed a mathematical framework to compare players, and in particular players spanning across different time periods in sports history. As illustrated in Figure 1, we aggregated cutoffs of historical ranking data into a single poset over top players by i) sampling from a data-driven random walk over the symmetric group and ii) adopting a notion of stochastic dominance. Ultimately, some pairs of players can be conclusively compared while others cannot. This approach departs from existing methods in that it relies on ranking data (rather than match-level data, which may or may not exist/be available) and formally comes to terms with the possibility that some comparisons are “too close to call.” Our methods leave ample room for further modeling choices, namely by adapting specifics of the constructions (3) and (4) as needed for the particular application in mind. In general, this works motivates the analytical study of data-driven sampling of permutations, be it with transpositions or other shuffling techniques. We anticipate this to be both challenging and of independent mathematical interest.
We implemented our methods using data from both the ATP and the WTA to find the GOATs in the respective categories. Our experimental results suggest that the “Big Three” of the ATP, that is Nadal, Federer, and Djokovic, are the ones that tend to come ahead as GOATs under most circumstances, whereas S. Williams and Graf are the ones that do so in the case of the WTA. Through our experiments, we also noted major differences in terms of career and dominance longevity for top players across the two associations.
Finally we remark that our methods, while initially motivated by this application in sports analytics, can be applied more broadly to other areas of operations management. For example, consider the area of route planning logistics, wherein many problems involve the frequent re-optimization of routes that start and end at fixed depots. In these settings, historical routes (e.g., optimized on a daily basis) can be compiled into a collection, in much in the same way as the historical rankings in Figure 1(a)—after all, a route is a permutation over locations to be visited. Therefore, methods similar to ours could produce a poset over locations, in this way “learning” (perhaps soft) precedence constraints from historical data. The distance between locations in terms of edges in Hasse diagram could intuitively encode the level of confidence in these constraints. Such information could be useful for the purpose of speeding-up integer programming solvers, through the addition of cuts that discard routes of poor quality. There are other examples of the usefulness of inferring precedence constraints, for example those stemming from the 2021 Amazon Last Mile Routing Research Challenge [merchan20242021]. Cook, Held, and Helsgaun [cook2024constrained] came first in the challenge by using historical last-mile delivery data to infer precedence constraints at a “zone” level (i.e., a higher level of aggregation than individual locations), in this way producing routes that closely-match those deemed of “high quality” by seasoned drivers who account for various practical aspects beyond merely route length. We therefore believe an extension of the ideas presented in this work could fall in line with recent calls  [winkenbach2024introduction] for “innovative solution methods” in route planning.